OpenStudy (anonymous):

Find the Inverse of f(x)= log(base 2) x-3

OpenStudy (jdoe0001):

I assume you've covered logarithms by now?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

well if you recall to find the "inverse relation" we simply swap about the variables, that is \(\bf f(x)={\color{brown}{ y}}=log_2({\color{blue}{ x}}-3)\qquad inverse\implies {\color{blue}{ x}}=log_2({\color{brown}{ y}}-3)\) then we just solve for "y" any ideas?

OpenStudy (anonymous):

Not sure

OpenStudy (jdoe0001):

\(\bf f(x)={\color{brown}{ y}}=log_2({\color{blue}{ x}}-3)\qquad inverse\implies {\color{blue}{ x}}=log_2({\color{brown}{ y}}-3) \\ \quad \\ \textit{log cancellation rule of }\large {\color{red}{ a}}^{log_{\color{red}{ a}}x}=x\qquad thus \\ \quad \\ x=log_2(y-3)\implies{ \large {\color{red}{ 2}}^x={\color{red}{ 2}}^{log_{\color{red}{ 2}}(y-3)}}\implies 2^x=y-3 \\ \quad \\ 2^x+3=y \iff f^{-1}(x)\)

OpenStudy (anonymous):

Ok, so the inverse is solving for Y?

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

well..yes

OpenStudy (jdoe0001):

yw